OpenStudy (sar9):

Jeremy is 7 years older than Jessica. In 9 years the sum of their ages will be 61. How old is Jeremy now?

OpenStudy (jing_wen):

\[(R-7)=S\] \[(R+9)+(S+9)=61\] \[R+S+18=61\] \[R+S=43\] since R-7=S R+(R-7)=43 2R=36 R=18 Jeremy is 18 years old

OpenStudy (sar9):

this was wrong

OpenStudy (mathmale):

Sar9: Have you read through Jing's discussion looking for anything that needs correction? You need to take more responsibility, and not just label someone else's help "wrong."

OpenStudy (sar9):

lol it was wrong though. dont comment if ur not gonna help.

OpenStudy (mathmale):

I see that jing has used the letter R to denote Jeremy's age, and the letter S to denote Jessica's age. Good going. I agree with this: "(R-7)=S)" Jeremy is older than Jessica.

OpenStudy (mathmale):

I also agree with (R+9)+(S+9)=61. sar9: Your turn. Solve this set of simultaneous equations for R and S.